# From Slow Bidirectional to Fast Causal Video Generators

### v1

@@ Abstract
Current video diffusion models achieve impressive generation quality but struggle in interactive applications due to bidirectional attention dependencies. The generation of a single frame requires the model to process the entire sequence, including the future. We address this limitation by adapting a pretrained bidirectional diffusion transformer to a causal transformer that generates frames on-the-fly. To further reduce latency, we extend distribution matching distillation (DMD) to videos, distilling 50-step diffusion model into a 4-step generator. To enable stable and high-quality distillation, we introduce a student initialization scheme based on teacher’s ODE trajectories, as well as an asymmetric distillation strategy that supervises a causal student model with a bidirectional teacher. This approach effectively mitigates error accumulation in autoregressive generation, allowing long-duration video synthesis despite training on short clips. Our model supports fast streaming generation of high quality videos at 9.4 FPS on a single GPU thanks to KV caching. Our approach also enables streaming video-to-video translation, image-to-video, and dynamic prompting in a zero-shot manner. We will release the code based on an open-source model in the future.
Figure 1: Traditional bidirectional diffusion models (top) deliver high-quality outputs but suffer from significant latency, taking 219 seconds to generate a 128-frame video. Users must wait for the entire sequence to complete before viewing any results. In contrast, we distill the bidirectional diffusion model into a few-step causal generator (bottom), dramatically reducing computational overhead. Our model (CausVid) achieves an initial latency of only 1.3 seconds, after which frames are generated continuously in a streaming fashion at approximately 9.4 FPS, facilitating interactive workflows for video content creation.
@@ 1 Introduction
In this paper, we introduce CausVid, a model designed for fast and interactive causal video generation. We design an autoregressive diffusion transformer architecture with causal dependencies between video frames. Similar to the popular decoder-only large language models (LLMs) [68, 6], our model achieves sample-efficient training by leveraging supervision from all input frames at each iteration, as well as efficient autoregressive inference through key-value (KV) caching. To further improve generation speed, we adapt distribution matching distillation (DMD) [104, 103], a few-step distillation approach originally designed for image diffusion models, to video data. Instead of naively distilling an autoregressive diffusion model [29, 8] into a few-step student, we propose an asymmetric distillation strategy where we distill the knowledge in a pretrained teacher diffusion model with bidirectional attention into our causal student model. We show that this asymmetric distillation approach significantly reduced error accumulation during autoregressive inference. This allows us to support autoregressively generating videos that are much longer than the ones seen during training. Comprehensive experiments demonstrate that our model achieves video quality on par with state-of-the-art bidirectional diffusion models while offering enhanced interactivity and speed. To our knowledge, this is the first causal video generation method that competes with bidirectional diffusion in terms of quality (Fig. 3 and Fig. 4). Additionally, we showcase the versatility of our method in tasks such as image-to-video generation, video-to-video translation, and dynamic prompting, all achievable with extremely low latency (Fig. 2).
@@ 5.1 Text to Video Generation
We evaluate the ability of our approach to generate short videos (5 to 10 seconds) and compare it against state-of-the-art methods: CogVideoX [100], OpenSORA [113], Pyramid Flow [29], and MovieGen [65]. As shown in Tab. 1, our method outperforms all baselines across all three key metrics: temporal quality, frame quality, and text alignment. Our model achieves the highest temporal quality score of $94.7$, indicating superior motion consistency and dynamic quality. In addition, our method shows notable improvements in frame quality and text alignment, scoring $64.4$ and $30.1$, respectively. We further report our method’s performance on the VBench-Long leaderboard in the supplementary material, where it achieves a total score of $82.85$, surpassing leading methods such as CogVideoX [100], OpenSORA [113], and Mira [30]. It also outperforms proprietary systems like Kling and Gen-3.
@@ 6 Discussion
Although our method is able to generate high-quality videos up to 30 seconds, we still observe quality degradation when generating videos that are extremely long. Developing more effective strategies to address error accumulation remains future work. Moreover, while the latency is significantly lower—by multiple orders of magnitude—compared to previous approaches, it remains constrained by the current VAE design, which necessitates the generation of five latent frames before producing any output pixels. Adopting a more efficient frame-wise VAE could reduce latency by an additional order of magnitude, significantly improving the model’s responsiveness. We also often notice discontinuities between consecutive video segments. Enhancing temporal smoothness is an important area for future work and may require improvements in both the generator and VAE design. In particular, we believe that incorporating temporal overlap when decoding consecutive video chunks, or introducing inter-block (causal) dependencies in the VAE, could alleviate this issue [62, 65]. Finally, while our method produces high-quality samples using the DMD objective, it comes with slightly reduced output diversity. This limitation is characteristic of reverse KL-based distribution matching approaches. Future work could explore alternative objectives such as EM-Distillation [96] and Score Implicit Matching [57], which may better preserve the diversity of outputs.
@@ A VBench-Long Leaderboard Results
We evaluate CausVid on the VBench-Long dataset using all 946 prompts across 16 standardized metrics. We refer readers to the VBench paper [27] for a detailed description of the metrics. As shown in Tab. 7, our method achieves state-of-the-art performance with the highest total score of 82.85. The radar plot in Fig. 8 visualizes our method’s comprehensive performance advantages. Our method is significantly ahead in several key metrics including dynamic degree, aesthetic quality, imaging quality, object class, multiple objects, and human action.
@@ C.1 Long-range Inconsistency
C.2 Temporal Discontinuity at Chunk Boundaries
Our model exhibits temporal discontinuity at the boundaries between adjacent chunks, as demonstrated by both visual examples on the website and the worse temporal flickering scores in Table 7. Some part of this issue stems from our VAE design, which processes single chunks of frames without incorporating inter-chunk dependencies, making it challenging to achieve smooth transitions at the boundaries. This limitation is further evidenced by our bidirectional teacher model, which, despite showing slight improvements, still exhibits similar temporal flickering issues due to the same VAE design (see Table 7).
Method Total Score Quality Score Semantic Score Subject Consistency Background Consistency Temporal Flickering Motion Smoothness Dynamic Degree Aesthetic Quality Imaging Quality Object Class Multiple Objects Human Action Color Spatial Relationship Scene Appearance Style Temporal Style Overall Consistency Vchitect 82.24 83.54 77.06 96.83 96.66 98.57 98.98 63.89 60.41 65.35 86.61 68.84 97.20 87.04 57.55 56.57 23.73 25.01 27.57 Jimeng 81.97 83.29 76.69 97.25 98.39 99.03 98.09 38.43 68.80 67.09 89.62 69.08 90.10 89.05 77.45 44.94 22.27 24.7 27.10 CogVideoX 81.61 82.75 77.04 96.23 96.52 98.66 96.92 70.97 61.98 62.90 85.23 62.11 99.40 82.81 66.35 53.20 24.91 25.38 27.59 Vidu 81.89 83.85 74.04 94.63 96.55 99.08 97.71 82.64 60.87 63.32 88.43 61.68 97.40 83.24 66.18 46.07 21.54 23.79 26.47 Kling 81.85 83.39 75.68 98.33 97.60 99.30 99.40 46.94 61.21 65.62 87.24 68.05 93.40 89.90 73.03 50.86 19.62 24.17 26.42 CogVideoX1.5-5B 82.17 82.78 79.76 96.87 97.35 98.88 98.31 50.93 62.79 65.02 87.47 69.65 97.20 87.55 80.25 52.91 24.89 25.19 27.30 Gen-3 82.32 84.11 75.17 97.10 96.62 98.61 99.23 60.14 63.34 66.82 87.81 53.64 96.40 80.90 65.09 54.57 24.31 24.71 26.69 CausVid (Ours) 82.85 83.71 79.40 93.33 96.65 93.85 97.64 82.20 66.49 70.20 93.70 74.68 99.80 82.37 66.72 55.86 24.13 25.23 27.52 Bidirectional Teacher 82.10 83.12 78.01 90.26 95.07 95.16 97.88 92.22 65.18 66.54 90.30 71.62 99.80 79.50 64.20 53.11 24.19 26.03 27.63

### v2

# From Slow Bidirectional to Fast Autoregressive Video Diffusion Models

@@ Abstract
Current video diffusion models achieve impressive generation quality but struggle in interactive applications due to bidirectional attention dependencies. The generation of a single frame requires the model to process the entire sequence, including the future. We address this limitation by adapting a pretrained bidirectional diffusion transformer to an autoregressive transformer that generates frames on-the-fly. To further reduce latency, we extend distribution matching distillation (DMD) to videos, distilling 50-step diffusion model into a 4-step generator. To enable stable and high-quality distillation, we introduce a student initialization scheme based on teacher’s ODE trajectories, as well as an asymmetric distillation strategy that supervises a causal student model with a bidirectional teacher. This approach effectively mitigates error accumulation in autoregressive generation, allowing long-duration video synthesis despite training on short clips. Our model achieves a total score of 84.27 on the VBench-Long benchmark, surpassing all previous video generation models. It enables fast streaming generation of high-quality videos at 9.4 FPS on a single GPU thanks to KV caching. Our approach also enables streaming video-to-video translation, image-to-video, and dynamic prompting in a zero-shot manner. We will release the code based on an open-source model in the future.
Figure 1: Traditional bidirectional diffusion models (top) deliver high-quality outputs but suffer from significant latency, taking 219 seconds to generate a 128-frame video. Users must wait for the entire sequence to complete before viewing any results. In contrast, we distill the bidirectional diffusion model into a few-step autoregressive generator (bottom), dramatically reducing computational overhead. Our model (CausVid) achieves an initial latency of only 1.3 seconds, after which frames are generated continuously in a streaming fashion at approximately 9.4 FPS, facilitating interactive workflows for video content creation.
@@ 1 Introduction
In this paper, we introduce CausVid, a model designed for fast and interactive causal video generation. We design an autoregressive diffusion transformer architecture with causal dependencies between video frames. Similar to the popular decoder-only large language models (LLMs) [66, 6], our model achieves sample-efficient training by leveraging supervision from all input frames at each iteration, as well as efficient autoregressive inference through key-value (KV) caching. To further improve generation speed, we adapt distribution matching distillation (DMD) [102, 101], a few-step distillation approach originally designed for image diffusion models, to video data. Instead of naively distilling an autoregressive diffusion model [29, 8] into a few-step student, we propose an asymmetric distillation strategy where we distill the knowledge in a pretrained teacher diffusion model with bidirectional attention into our causal student model. We show that this asymmetric distillation approach significantly reduced error accumulation during autoregressive inference. This allows us to support autoregressively generating videos that are much longer than the ones seen during training. Comprehensive experiments demonstrate that our model achieves video quality on par with state-of-the-art bidirectional diffusion models while offering enhanced interactivity and speed. To our knowledge, this is the first autoregressive video generation method that competes with bidirectional diffusion in terms of quality (Fig. 3 and Fig. 4). Additionally, we showcase the versatility of our method in tasks such as image-to-video generation, video-to-video translation, and dynamic prompting, all achievable with extremely low latency (Fig. 2).
@@ 5.1 Text to Video Generation
We evaluate the ability of our approach to generate short videos (5 to 10 seconds) and compare it against state-of-the-art methods: CogVideoX [98], OpenSORA [111], Pyramid Flow [29], and MovieGen [63]. As shown in Tab. 1, our method outperforms all baselines across all three key metrics: temporal quality, frame quality, and text alignment. Our model achieves the highest temporal quality score of $94.7$, indicating superior motion consistency and dynamic quality. In addition, our method shows notable improvements in frame quality and text alignment, scoring $64.4$ and $30.1$, respectively. In the supplementary material, we present our method’s performance on the VBench-Long leaderboard, achieving a total score of $84.27$ and securing first place among all officially evaluated video generation models.
@@ 6 Discussion
Although our method is able to generate high-quality videos up to 30 seconds, we still observe quality degradation when generating videos that are extremely long. Developing more effective strategies to address error accumulation remains future work. Moreover, while the latency is significantly lower—by multiple orders of magnitude—compared to previous approaches, it remains constrained by the current VAE design, which necessitates the generation of five latent frames before producing any output pixels. Adopting a more efficient frame-wise VAE could reduce latency by an additional order of magnitude, significantly improving the model’s responsiveness. Finally, while our method produces high-quality samples using the DMD objective, it comes with reduced output diversity. This limitation is characteristic of reverse KL-based distribution matching approaches. Future work could explore alternative objectives such as EM-Distillation [94] and Score Implicit Matching [56], which may better preserve the diversity of outputs.
@@ A VBench-Long Leaderboard Results
We evaluate CausVid on the VBench-Long dataset using all 946 prompts across 16 standardized metrics. We refer readers to the VBench paper [27] for a detailed description of the metrics. As shown in Tab. 7, our method achieves state-of-the-art performance with the highest total score of 84.27. The radar plot in Fig. 8 visualizes our method’s comprehensive performance advantages. Our method is significantly ahead in several key metrics including dynamic degree, aesthetic quality, imaging quality, object class, multiple objects, and human action. More details can be found on the official benchmark website (https://huggingface.co/spaces/Vchitect/VBench_Leaderboard).
@@ C.1 Long-range Inconsistency
Method Total Score Quality Score Semantic Score Subject Consistency Background Consistency Temporal Flickering Motion Smoothness Dynamic Degree Aesthetic Quality Imaging Quality Object Class Multiple Objects Human Action Color Spatial Relationship Scene Appearance Style Temporal Style Overall Consistency Vchitect 82.24 83.54 77.06 96.83 96.66 98.57 98.98 63.89 60.41 65.35 86.61 68.84 97.20 87.04 57.55 56.57 23.73 25.01 27.57 Jimeng 81.97 83.29 76.69 97.25 98.39 99.03 98.09 38.43 68.80 67.09 89.62 69.08 90.10 89.05 77.45 44.94 22.27 24.7 27.10 CogVideoX 81.61 82.75 77.04 96.23 96.52 98.66 96.92 70.97 61.98 62.90 85.23 62.11 99.40 82.81 66.35 53.20 24.91 25.38 27.59 Vidu 81.89 83.85 74.04 94.63 96.55 99.08 97.71 82.64 60.87 63.32 88.43 61.68 97.40 83.24 66.18 46.07 21.54 23.79 26.47 Kling 81.85 83.39 75.68 98.33 97.60 99.30 99.40 46.94 61.21 65.62 87.24 68.05 93.40 89.90 73.03 50.86 19.62 24.17 26.42 CogVideoX1.5-5B 82.17 82.78 79.76 96.87 97.35 98.88 98.31 50.93 62.79 65.02 87.47 69.65 97.20 87.55 80.25 52.91 24.89 25.19 27.30 Gen-3 82.32 84.11 75.17 97.10 96.62 98.61 99.23 60.14 63.34 66.82 87.81 53.64 96.40 80.90 65.09 54.57 24.31 24.71 26.69 CausVid (Ours) 84.27 85.65 78.75 97.53 97.19 96.24 98.05 92.69 64.15 68.88 92.99 72.15 99.80 80.17 64.65 56.58 24.27 25.33 27.51